If 20 mL of 1.0M HCl reacts with 0.050g of solid magnesium. Calculate the moles of each product formed. Be sure to explicitly show how the limiting reactant was determined.

The balanced equation for this reaction is as follows

Mg(s)+2HCl(aq)---->MgCl2(aq)+H2(g)
Answer #1


Mg(s)+2HCl(aq)---->MgCl2(aq)+H2(g)


No of mol of Mg = 0.05/24 = 0.0021 mol

No of mol of HCl = 20/1000*1 = 0.02 mol

limiting reactant = Mg

no of mol of MgCl2 produced = 0.0021 mol

no of mol of H2 produced = 0.0021 mol